Marchete Company produces a single product. They have recently received the results of a market survey that indicates that they can increase the retail price of their product by 10% without losing customers or market share. All other costs will remain unchanged. Their most recent CVP analysis is presented below.

Current
Units sold   950
Sales Price per Unit   $120
Variable Cost per Unit   $99
Contribution Margin per Unit   $21
Fixed Costs   $17,829
Break-Even (in units)   849
Break-Even (in dollars)   $101,880
Sales   $114,000
Variable Costs   $94,050
Contribution Margin   $19,950
Fixed Costs   $17,829
Net Income (loss)   $2,121
If they enact the 10% price increase, what will be their new break-even point in units and dollars? If required, round final answers to nearest whole number.

New Price
Break-even (in units)  
???
Break-even (in dollars)   $
???
Answer #1

Contribution margin per unit after price increase = 120*110% - 99

= $33 per unit

Break even units = Fixed costs/Contribution margin per unit

= 17829/33

= 540.27 units

CM Ratio = 33/132 = 25%

Hence, break even dollars = Fixed costs/CM Ratio

= 17829/25%